OpenStudy (anonymous):

Solve for x: 5 ( x - 2 ) ≤ 30

OpenStudy (kuoministers):

first step is to expand the brackets

OpenStudy (kuoministers):

so i hope you know how to do that

OpenStudy (kuoministers):

e.g 3(4 + x) 12 + 3x

OpenStudy (anonymous):

yes

OpenStudy (kuoministers):

so can you expand the brackets? 5(x - 2)

OpenStudy (anonymous):

5x -2 ≤ 30

OpenStudy (anonymous):

sry--

OpenStudy (anonymous):

I would just divide the five out since it is a factor > 0

OpenStudy (anonymous):

5x - 10 ≤ 30

OpenStudy (kuoministers):

OK GOOD now we need to add 10 to get rid of -10 right? so we add 10 to both sides

OpenStudy (anonymous):

x-2 <= 6 and then isolate x by adding 2 to both sides x <= 8

OpenStudy (kuoministers):

http://openstudy.com/code-of-conduct Never give away the answer!!!

OpenStudy (anonymous):

okay, sry

OpenStudy (anonymous):

i can still process through

OpenStudy (anonymous):

5x > 20

OpenStudy (kuoministers):

well since answer is already given 5x - 10 = 30 we would add 10 not minus 5x = 40

OpenStudy (anonymous):

okay-- 40/5 = 8

OpenStudy (kuoministers):

i meant 5x ≤ 30 btw

OpenStudy (kuoministers):

so x ≤ 8

OpenStudy (kuoministers):

yes gw

OpenStudy (kuoministers):

If you think my response helped you please click the best response box

OpenStudy (anonymous):

absolutely! maybe back in a bit thanx for your help